INAF LBC+MODS queue observing, 2011 November 27-28

Observer: K. Boutsia, M. Fumana
Telescope Operator: D. Gonzales-Huerta
Support Astronomer: O. Kuhn
Instrument: LBC+MODS

Summary

Clear sky at the beginning of the night, probably photometric . The seeing has been between 0.6'' and 0.9'' for the whole night.
We started with LBC and finished Clementini and Giallongo (CL0024).
LBC blue was not able to hold the focus at the beginning of the night - elongated images - and this was probably caused by the temperature gradient.
LBC red also gave problems with the focus. We actually had to interrupt 2 OBs becuase the stars wer very comatic on chip1 - there was a gradient in the psf - comatic on the right side of the CCD.
Then we passed to MODS, we finished Donnarumma and started Nonino macs0717.
30min lost to acquire a MOS mask for a second time because of the telescope unwrapping.
At the end of the night we had an error with MODS, that it could not lock the laser on the right side ... we were at lowish elevation (45).
We noticed that after the MOS acquisition with MODS, the stars were still not well centered in the slits, actually the stars on the upper side were better centered than the ones on the lower side. We had to adjust manually. We also noticed in a mask where some spectra were visible, that they were all on the upper side of the detector. Does this mean that the mask was mis-aligned?? All the bright reference sources were within their silts and we could see the spectra...

Data Summary

Project Objects Filter ExpTime Notes
Giallongo CL0024+16 g-sloan - r-sloan 2hr30min this program is considered completed
Clementini AndXXV Bbess - Vbess 20min seeing ~0.8'' but on the blue side the images are elongated
Clementini AndXXVII Bbess - Vbess 20min seeing ~0.8'' but on the blue side the images are elongated
STD SA113 Bbess / g-sloan - Vbess / r-sloan   4 positions -airmass 1.2
STD Stetson_L92 Bbess / g-sloan - Vbess / r-sloan   4 positions - airmass 1.46
Donnarruma MAC0717b1 MODS red grating 2250sec STD 5''slit + arc +flat +bias (for calib files look a UT 20111126)
Donnarruma MAC0717b2 MODS red grating 2250sec STD 5''slit + arc +flat +bias (for calib files look a UT 20111126)
Nonino macsj0717+3745 MODS dual grating 2hr30min STD 5''slit + arc +flat +bias (for calib files look a UT 20111126)

Details

00:10 Turn ON LBC

00:15 Take some test Bias frames to see if everything is working
  • 25Bias_Bino - the lights are OFF but vent doors are open

00:33 Taking SkyFlats

00:55 Slew to STD SA113 - dofpia
  • dofpia,/First
 -486     64     57  -1638   -221   -445    -10   0.94   0.82     red
-9934      0      0      0      0      0      0   0.00   3.92    blue

 -1384    219     49    166   -100    184    100   1.07   0.57     red
 -1002   -173   -405  -1575   -319   -330    -17   1.75   0.86    blue

 -595    176     99   -499    -51    244     67   1.18   0.37     red
 -360   -408    139   -232    -86    119    100   1.96   0.25    blue

 306     30    133     29    -37    101     53   1.07   0.16     red
  611   -381   -146    425    -90   -243     57   1.86   0.37    blue

01:06 SA113_Bg_Vr
  • 4 positions in filters Bbess-g-SLOAB and Vbess-r-SLOAN
  • airmass 1.2
  • g band seemed saturated - repeat

01:20 Slew to Stetson_L92_Bg_Vr
  • 4 positions in filters Bbess-g-SLOAB and Vbess-r-SLOAN
  • airmass 1.46

01:28 Slew to Pointing Star ACT0364 for collimation and pointing correction
  • dofpia
4094   -150     -7    535   -711   -143     13   1.09   1.67     red
 1647     -5    159   -428   -587   -165     45   1.71   0.74    blue

166    -23    -98   -206     39   -174     35   0.94   0.18     red
 38   -170    -43     11      7   -152     65   1.12   0.14    blue

  • dimm seeing 0.8-0.9

01:38 M31_AndXXV_400s_2
  • dofpia
293    -45    -68     90    100   -193     21   0.93   0.20     red
1082    -86   -176   -190    211   -131     17   1.25   0.46    blue

-1340    -62     54    157   -670    350     15   0.95   0.66     red
543    178     40    193     11   -179     20   1.27   0.27    blue

1663    -89    -75    -57    527   -302      0   0.98   0.73     red
225    -48     85     56     78    -78     15   1.34   0.14    blue

-46    -52    -65    -53    201   -127     20   1.03   0.15     red
248    -54    112    -55     40    -67     -4   1.27   0.14    blue
* dimm seeing 0.8

01:45 M31_AndXXV_400s_2
  • dimm seeing 0.8
  • seeing measured on the images: Blue 1.0''/ Red 0.8''
  • blue channel are elongated on 2nd image
  • seeing measured on the images: Blue 1.1''/ Red 0.8''
  • blue doesn't send correction, red does
  • dimm seeing 0.8
  • 3th blue is elongated

02:09 Slew M31_AndXXVII_400s_2
  • dofpia
1053    -56     45    180   -233   -212     -8   1.00   0.46     red
3514    -74   -198    279     61   -210    -34   1.46   1.40    blue

297    -42    -57    -90    208   -167      6   0.97   0.20     red
-2     66    114     67   -154      0    -10   1.53   0.12    blue

02:12 M31_AndXXVII_400s_2
  • dimm seeing 0.9
  • seeing measured on the images: Blue 0.97''/ Red 0.72''
  • seeing measured on the images: Blue 1.0''/ Red 0.8''

02:35 Run focus OB
  • dofpia and rangebal on a target in Giallongo_CL0024+16_GR1 field
 1127     -3    -29   -421    261   -217    -51   0.96   0.52     red
2234    155    -57   -106    397    -74    -29   1.69   0.91    blue

-144    109    194    310     43    -71    -14   0.98   0.19     red
-164    174     43    112    -12     88    -13   1.60   0.14    blue

02:45 Giallongo_CL0024+16_GR4
  • dimm seeing 0.8
  • seeing measured on the images: Blue 1.0''/ Red 0.8''
  • seeing measured on the images: Blue 0.9''/ Red 0.8''
  • during the second half of the OB the blue images start being elongated - it does not hold focus on the blue side
  • seeing measured on the images: Blue 0.8''/ Red 0.7''

03:25 dofpia
 1311     -4   -128    132    216    -91    -67   0.94   0.54     red
 1359     94   -205    -83   -385    -64    -50   1.45   0.57    blue

-161     88   -100   -208    109     93    -36   0.94   0.16     red
 337     45   -157      0    -52    -53    -52   1.37   0.17    blue

03:28 Giallongo_CL0024+16_GR5
  • seeing measured on the images: Blue 0.8''/ Red 0.7''
  • seeing measured on the images: Blue 0.85''/ Red 0.7''

04:07 dofpia
-1034     43    -43     58    -72    -43    -30   0.67   0.42     red
 501   -189   -102    160     91      3    -26   1.16   0.24    blue

  47    105     67    -24    -73    -44    -30   0.66   0.10     red
 623    114    256     -8     84    -13    -26   1.02   0.27    blue

  -274    -22     52     53    -52     26    -15   0.63   0.14     red
 -666    -58   -108     22     89    112     -5   0.94   0.29    blue
 
  -6     29    -56     46     24    -14    -22   0.59   0.08     red
 180   -116     88    119     75     54      6   1.06   0.14    blue

04:14 Giallongo_CL0024+16_GR6
  • seeing measured on the images: ~0.6''
  • in the red images the sources look like donuts
  • we stop the OB almost half way (4blue images + 7red) because the last red images are not useful

04:34 dofpia
-1798    -66     35   -128     57   -125      5   0.39   0.72     red
 519    -69   -101    -85     60   -186     -5   1.03   0.26    blue

 -282    -40      6    131     -4    -24      2   0.45   0.15     red
  200   -112     30      7     -4    -83    -24   1.02   0.13    blue

 -164    -13     36   -106    -39     12      3   0.47   0.11     red
  -247    -93     69      0    -54     80      8   0.96   0.14    blue

04:38 Giallongo_CL0024+16_GR7
  • seeing measured on the images: ~0.6''
  • dimm seeing 0.7''
  • many sources on the red side look saturated... (?)

05:18 dofpia
 -113    -15    -56    -68    448   -134     -9   0.52   0.24     red
 -552   -159   -151    216    124    -15      8   1.25   0.27    blue

   9     28     41     14    -53     -5      3   0.53   0.08     red
  235    164     88    -97     49   -100    -12   1.30   0.16    blue

05:22 Giallongo_CL0024+16_GR8
  • seeing measured on the images: ~0.6''
  • dimm seeing 1.0
  • dimm seeing 0.7
  • seeing measured on the images: ~0.6''

06:02 dofpia
498     59    109     52    -31   -119     -6   0.64   0.23     red
4    211     12     52    -57     -7     -3   1.15   0.11    blue

 172     30    -50    -38     29    -69     -1   0.69   0.12     red
-189    -89    -91   -142     66    116      2   1.22   0.15    blue

06:05 Giallongo_CL0024+16_GR4
  • dimm seeing 1.0
  • seeing measured on the images: 0.6''-0.7''
  • interrupted at the middle again because on the red side there is a gradient in the psf - on the right side the sources are comatic (9red + 6 blue) - 18min

06:30 SWITCH TO MODS

06:58 acqMODS mods.7189.201130b2.acq 0001 - 0005
  • dimm seeing 0.9'' - gcs 0.9''
  • mods1r.20111128.0001.fits - slit image
  • mods1r.20111128.0002.fits - field image
  • mods1r.20111128.0003.fits - throughslit image
  • we did a bit of manual adjustment because the stars at the lower part seemed to be a bit offset (-0.246arcsec in dY)

07:25 execMODS mods.7189.201130b2_1exp.obs 0006
  • guider seeing <1.0''

08:06 acqMODS mods.7190.201130b1.acq 0007-0010
  • mods1r.20111128.0007.fits - slit image
  • mods1r.20111128.0008.fits - field image
  • mods1r.20111128.0009.fits - throughslit image
  • we did a bit of manual adjustment because the stars at the lower part seemed to be a bit offset (-0.5 arcsec in dY)

08:25 execMODS mods.7190.201130b1_1exp.obs 0011
  • guider seeing <0.9''

09:06 acqMODS macsj0717_lbt_mods_27_IX_2011.acq 0012-0015
  • mods1r.20111128.0012.fits - slit image
  • mods1r.20111128.0013.fits - field image
  • mods1r.20111128.0014.fits - throughslit image
  • we did a bit of manual adjustment because the stars at the lower part seemed to be a bit offset (-0.74 arcsec in dY)

09:25 macsj0717_lbt_mods_27_IX_2011_exp16.obs mods1r 0016-0017 / mods1b 0001-0002
  • dimm seeing 0.6''

10:06 presets cancelled - we crossed the meridian and the telescope is unwrapping
  • we need to repeat the acquisition

10:08 acqMODS macsj0717_lbt_mods_27_IX_2011.acq 0018-0019

10:19 we had to abort the preset again because of azimuth unwrapping problem again
  • probably before the preset was cancelled for another reason

10:24 acqMODS macsj0717_lbt_mods_27_IX_2011.acq 0020-0022

10:35 execMODS macsj0717_lbt_mods_27_IX_2011_exp16.obs mods1r 0023-0030 / mods1b 0003-0010
  • dimm seeing 0.86''
  • abort because it is getting bright - 8 exposures in each channel (10 total)

12:53 acqMODS standard_star.acq 0031 - 0032

13:04 execMODS standard_star.obs 0033
  • 5arcsec slit

13:10 execMODS standard_star.obs mods1r 0034 / mods1b 0011
  • repeat with dual grating
  • the OBJECT name on the blue side is "macsj0717+3745" and not "BD+75d325 dual grating" as expected, because we did not update the script
  • 5arcsec slit
>> imcslock
** ERROR: IMCSLOCK RedIMCS=Fail BlueIMCS=LockedOn
** Abort, Retry or Ignore? > retry

** standard_star.obs aborted with errors at line 8 command 'imcslock'
  • we got the same error again - but this time we ignore and take the standard with the blue grating also
>> imcslock
** ERROR: IMCSLOCK RedIMCS=Fail BlueIMCS=LockedOn
** Abort, Retry or Ignore? > i
** Ignoring error and continuing...

13:18 acqMODS standard_star1arcsec.acq 0035-0036
  • mods1r.20111128.0035.fits - slit image
  • we lost the guide star - we calculated the offsets and we try to observe without guiding

13:26 execMODS standard_star1arcsec.obs mods1r 0034 / mods1b 0012
  • the lazer gave error again for the red side - we take the exposure anyway but it might not be centered anymore because we are not guiding
  • no star in the slit .... no standard with 1'' slit

13:30 END OF NIGHT

  • Seeing reported by GCS during MODS observatioins:
    20111128fwhm.jpg

-- KonstantinaBoutsia - 28 Nov 2011
Topic revision: r8 - 23 Feb 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback